# Unitarization of the Sommerfeld enhancement through the renormalization group

### v1

@@ 3.1 Origin of the unitarity violation
An analogous situation is expected for the Sommerfeld enhancement. Let us therefore consider what the actual unitarized behavior of the Sommerfeld enhancement should be, in light of the mechanisms that regulate singularities in simpler contexts such as the $s$-channel resonance.131313 In the presence of a shallow bound state, if we describe the self-scattering amplitude using an effective field theory, we must introduce not only fields for the scattering particles but also an additional field corresponding to the bound state that couples to them. In this framework, the self-scattering process is described as an $s$-channel resonance mediated by the bound-state field.For example, in the zero-range approximation of the $s$-wave elastic scattering amplitude in the absence of annihilation, namely considering only a positive scattering length $a_{0}>0$ in Eq. (2.18), the self-scattering amplitude is given by
@@
which has a bound state pole at $p_{B}=i/a_{0}$. When the annihilation effect is turned on, the scattering length acquires an imaginary part, and the bound-state pole shifts off the imaginary axis, signaling its instability. Specifically, if the inverse scattering length shifts as $1/a\to-1/a-i\gamma$ with $\gamma>0$ [^14] The positivity of the decay width of the bound state $\Gamma=2\gamma/\mu a_{0}$ requires that the sign of $\gamma$ be positive., the optical theorem implies that the annihilation cross section becomes [35, 36]
@@ 3.4 Behavior of the improved scattering amplitude
This observation provides a simple prescription for the improved Sommerfeld enhancement in the small-momentum regime, especially, for the $s$-wave case, as originally proposed in [9]. Neglecting the real part of $f^{S}(p)$, the improved Jost function can be approximated as $\mathscr{J}_{\ell}^{\rm imp}(p)\simeq\alpha_{0}+i\gamma_{0}(p+p_{c})\simeq\mathscr{J}_{\ell}(p+p_{c})$, with $p_{c}$ being a regularising momentum given by $p_{c}={\rm Im}f_{\ell}^{S}(p)/\gamma_{0}^{2}=\mu(\sigma_{0}v)^{\rm ann,0}/4\pi\gamma_{0}^{2}$. Therefore, the improved Sommerfeld factor can also be approximated by the naive one evaluated at a momentum shifted by the critical value, $S^{\rm imp}_{0}(p)\simeq S_{0}(p+p_{c})$.
@@
Fig. 2 shows the Sommerfeld factors as functions of momentum $p$ (normalized by the well width $R$) for the $s$-wave (left panel) and the $\ell$-wave (right panel). We set the mass parameter and the coupling constant to $\mu R=10$ and $\alpha_{D}=10^{-2}$, respectively. In both panels, the dashed and solid lines of the same color form a pair: the dashed line represents the conventional Sommerfeld factor $S_{\ell}(p)$ given in Eq. (2.25), while the corresponding solid line indicates the RG-improved Sommerfeld factor $S_{\ell}^{\rm imp}(p)$ in Eq. (3.21). The red solid line shows the ratio of the unitarity-bound annihilation cross section, which is given by the RHS of Eq. (3.3), to the tree-level annihilation cross section Eq. (3.33). For the $s$-wave case, we plot the Sommerfeld factors with the well depth chosen near the first zero-energy resonance, at $p_{V}=\pi/2-10^{-3}$ (purple), $p_{V}=\pi/2-10^{-5}$ (blue), and near the second zero-energy resonance, at $p_{V}=3\pi/2-10^{-5}$ (green). For the purple line, which is sufficiently far from the zero-energy resonance, the conventional Sommerfeld factor remains well below the unitarity bound and agrees well with the improved one. In contrast, for the blue and green lines, which are closer to the zero-energy resonance, both show unphysical behavior at small momenta, exceeding the unitarity bound. The RG-improved Sommerfeld factor resolves these issues and yields results consistent with unitarity. For the $p$-wave case, we plot the points $p_{V}=\pi-10^{-3}$ (purple) and $p_{V}=\pi-10^{-5}$ (blue) near the first zero-energy resonance, and $p_{V}=3\pi-10^{-5}$ (green) near the third zero-energy resonance. Unlike the $s$-wave case, the conventional Sommerfeld factors for all three parameter choices exhibit spike-like structures [8]. These arise from the fact that the $p$-wave can support resonant states, corresponding to zeros of the Jost function in the region of positive momentum. In particular, the blue and green lines exhibit resonant enhancements that exceed the unitarity bound. In contrast, the RG-improved Sommerfeld factor, shown as dashed lines, resolves such unphysical behavior: while it coincides with the conventional one for the purple line, which does not violate unitarity, the resonance peaks in the blue and green lines are properly regulated, yielding results consistent with unitarity.
@@ 4 Discussion
In this section, we provide further insights into the results obtained in the previous section by addressing two complementary aspects. First, we revisit the RG improvement procedure from the perspective of Wilosnian renormalization group [40], clarifying the validity and limitations of our method. This analysis helps identify conditions under which the RG improvement may fail or require modification. Second, as an advanced topic concerning the Sommerfeld effect, we briefly discuss the bound-state formation process [18, 41, 42], which can exhibit behaviors that violate the unitarity bound and therefore require careful treatment. These discussions aim to deepen the understanding of our formalism and its potential applicability, though they remain at a qualitative level and a more detailed treatment is left for future research.
@@ 4.1 Insights from Wilsonian renormalization group
where $G_{L}$ is a Green function aasociated with the long-range potential $V_{L}$. (For its explicit form, see [43]; however, the details are not essential for the following discussion.) The quantity $V_{S}(\Lambda)$ can be interpreted as the effective potential at the scale $p\sim\Lambda$, and the renormalization group equation describes its flow as a function of the momentum scale. The fixed points of the renormalization group equation (4.1) and their stability in the limit $p\to 0$ characterize the nature of the short-range contributions to the scattering process.
@@
where $\tilde{\delta}^{S}_{\ell}(p)$ is defined by the difference between the total phase shift $\delta_{\ell}(p)$ and that due to the pure long-range potential $\delta^{L}_{\ell}(p)$. $\mathcal{M}_{\ell}(p)$ is given by $\mathcal{M}_{\ell}(p)=\lim_{r\to 0}((pr)^{\ell}\chi^{+}_{\ell,p}(r))/2^{\ell}\ell!\mathscr{J}_{\ell}(p)$, which is constructed using the Jost solution $\chi^{+}_{\ell,p}(r)$ (the solution of the Schrödinger equation (2.10) with boundary conditions imposed at infinity, $\chi^{+}_{\ell,p}(r)/h^{+}_{\ell}(pr)\to 1,r\to\infty$). The extra factors appearing in the modified effective range expansion Eq. (4.2) compared to the standard one Eq. (2.18) serve to remove the non-analytic momentum dependence arising from the long-range potential. This expansion has been used to subtract long-range effects such as the Coulomb force or one-pion exchange between nucleons [45, 29], thereby enabling the extraction of low-energy properties of the strong interaction. Therefore, if one allows the shape parameters $\tilde{a}_{\ell},\tilde{r}_{e}$ to take complex values to accounts for annihilation processes, Eq. (4.2) is expected to provide a useful parametrization of the self-scattering amplitude in situations where the short-range effect is nonperturbative, though it remains applicable even when the short-range effect is perturbative.
@@
Once the potential is regularized by some method and the Schrödinger equation is solved, the resulting $S$-matrix must reduce to the form of the modified effective range expansion, as it provides a general parametrization of the scattering amplitude for low-energy regime in the case where the potential comprises two separate contributions. In this case, the shape parameters become functions of the regularization parameters (and parameres contained in the long-range potential as well), but they are fixed to definite numerical values by matching to physical observables—namely, through renormalization. This approach is, in essence, equivalent to using the modified effective range expansion and fitting its parameters to scattering data at a suitable momentum scale. The key point, then, is whether the effect of the short-range potential is perturbative or nonperturbative. If the short-range interaction is perturbative, the corresponding contribution to the scattering amplitude can be computed directly from the UV theory using perturbation theory. In such cases, the RG–based treatment we have proposed is expected to provide a sufficiently accurate and much simpler approximation, particularly for computing unitarized Sommerfeld factors, compared to performing a matching based on the modified effective range expansion. Indeed, Ref. [9] discusses that, for the $s$-wave case, solving the Schrödinger equation with the short-range effect included yields results that agree well numerically with our expression given in Sec. 3.3 when the underlying short-range effect is perturbative. Establishing a more general correspondence between the modified effective range expansion and the RG-based approach in cases where the short-range effect is perturbative for general partial waves remains a future work.
@@ 4.2 Bound state formation
In the preceding discussion, we have considered only the simple process of direct annihilation. As a higher-order contribution to the annihilation process, it is known that slowly moving particles can form bound states through the emission of soft bosons that mediate their long-range interactions, which is referred to as bound state formation (BF) [18, 41, 42]. It has been pointed out that BF also exhibits resonance peaks at parameter values where the self-scattering cross section and the Sommerfeld factor resonate, in contradiction with the partial-wave unitarity bound [46, 47].
This result can be understood through the following simple consideration. For simplicity, let us assume that the mediator particle is a scalar boson $\varphi$, and that the transition from the scattering state $\mathcal{S}$ to a bound state $\mathcal{B}$ proceeds via the emission of a light boson with momentum $k$, mediated by a scalar transition operator $V_{k}$. At low energies, the initial state is dominated by the $s$-wave contribution and then, under the leading-order DWBA, the matrix element for this process is roughly given by
@@ 5 Conclusion
In the discussion, we examined our approach from the viewpoint of the Wilsonian renormalization group. From this perspective, our method under the assumption of a perturbative UV theory can be understood as a perturbative expansion around the Gaussian fixed point. The improvement of the amplitude through perturbative renormalization shares a parallel structure with the standard renormalized perturbation theory applied to weakly coupled field theories such as QED. Hence, the renormalization group approach is expected to offer a more tractable framework for unitarizing the Sommerfeld enhancement in systems with weak short-range interactions, compared to approaches based on exact solutions to the Schrödinger equation. A precise correspondence between these two frameworks for arbitrary partial waves remains an open question.
@@ A.1 Riccati-Bessel functions
$j_{\ell}(z)=z^{\ell+1}\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!(2\ell+2n+1)!!}\left(\frac{z^{2}}{2}\right)^{n},\quad n_{\ell}(z)=\frac{1}{z^{\ell}}\sum^{\infty}_{n=0}\frac{(-1)^{n}(2\ell-2-1)!!}{n!}\left(\frac{z^{2}}{2}\right)^{n}.$ (A.2)
@@
$h^{\pm}_{0}(z)=e^{\pm z},$ $\quad h_{1}^{\pm}(z)=\left(1+\frac{i}{z}\right)e^{\pm i(z-\pi/2)}.$ (A.9)

### v2

@@ 3.1 Origin of the unitarity violation
An analogous situation is expected for the Sommerfeld enhancement. Let us therefore consider what the actual unitarized behavior of the Sommerfeld enhancement should be, in light of the mechanisms that regulate singularities in simpler contexts such as the $s$-channel resonance. [^13] In the presence of a shallow bound state, if we describe the self-scattering amplitude using an effective field theory, we must introduce not only fields for the scattering particles but also an additional field corresponding to the bound state that couples to them. In this framework, the self-scattering process is described as an $s$-channel resonance mediated by the bound-state field.For example, in the zero-range approximation [36] of the $s$-wave elastic scattering amplitude in the absence of annihilation, namely considering only a positive scattering length $a_{0}>0$ in Eq. (2.18), the self-scattering amplitude is given by
@@
which has a bound state pole at $p_{B}=i/a_{0}$. When the annihilation effect is turned on, the scattering length acquires an imaginary part, and the bound-state pole shifts off the imaginary axis, signaling its instability. Specifically, if the inverse scattering length shifts as $1/a\to-1/a-i\gamma$ with $\gamma>0$ [^14] The positivity of the decay width of the bound state $\Gamma=2\gamma/\mu a_{0}$ requires that the sign of $\gamma$ be positive., the optical theorem implies that the annihilation cross section becomes [36, 37, 38]
@@ 3.3 Unitarization of Sommerfeld enhancement via the RG
Let us remark that the simplified version of the amplitude Eq. (3.20) obtained using the RG method can also be derived on the basis of the following, more intuitive argument. In general, solving the Schrödinger equation that includes a highly singular short-range potential—such as one given by a delta function and its derivatives—is not straightforward (see [9, 10] and the discussion in Sec 4.1). Suppose, however, that it can be solved in some sense (for example, under an appropriate regularization and matching). Then, as in ordinary potential scattering, the Schrödinger equation admits scattering solution, $\psi_{\ell,p}^{\rm mod}(r)$, and its behavior near the origin should be expressed in terms of a Jost function $\mathscr{J}_{\ell}^{\rm mod}(p)$ that is modified to incorporate the effects of the singular potential, $\psi^{\rm mod}_{\ell,p}(r)\to j_{\ell}(pr)/\mathscr{J}_{\ell}^{\rm mod}(p)$. Combining this with the two-potential formula Eq. (2.21), the modified self-scattering amplitude $f_{\ell}^{\rm mod}(p)$ including the effects of the short-range interaction becomes
$f_{\ell}^{\rm mod}(p)=f^{L}_{\ell}(p)+\frac{1}{\mathscr{J}_{\ell}(p)}\frac{1}{\mathscr{J}_{\ell}^{\rm mod}(p)}f_{\ell}^{S}(p).$ (3.21)
On the other hand, taking into account the role of the Jost function in scattering theory, the modified self-scattering amplitude should be written using the modified Jost function as
$f_{\ell}^{\rm mod}(p)=\frac{\mathscr{J}^{\rm mod}_{\ell}(-p)-\mathscr{J}_{\ell}^{\rm mod}(p)}{2ip\mathscr{J}^{\rm mod}_{\ell}(p)}.$ (3.22)
Only the part of the modified Jost function that is odd in $p$ contributes to the numerator above. From Eq. (2.17), the source of the $p$-odd piece is the unitarity cut arising from self-scattering due to the long-range potential. Therefore, as a first approximation, we may take all contributions of the short-range potential—including annihilation effects—to the modified Jost function to be $p$-even, and thus the $\mathscr{J}^{\rm mod}_{\ell}(p)$ in the numerator can be replaced by the Jost function $\mathscr{J}_{\ell}(p)$ due solely to the long-range potential. Then we find the self-consistent equation for the modified Jost function
$\frac{\mathscr{J}_{\ell}(-p)-\mathscr{J}_{\ell}(p)}{2ip\mathscr{J}^{\rm mod}_{\ell}(p)}=f^{L}_{\ell}(p)+\frac{1}{\mathscr{J}_{\ell}(p)}\frac{1}{\mathscr{J}_{\ell}^{\rm mod}(p)}f_{\ell}^{S}(p),$ (3.23)
whose solution leads us the improved Jost function in Eq. (3.20). This line of reasoning may also provide an intuitive understanding of when Eq. (3.20) obtained from the RG method breaks down: if the contributions of the long-range and short-range potentials to the scattering become comparable, the effect of the unitarity cut can no longer be represented by the long-range Jost function alone, and the numerator-replacement trick ceases to be applicable.
@@ 3.4 Behavior of the improved scattering amplitude
This observation provides a simple prescription for the improved Sommerfeld enhancement in the small-momentum regime, especially, for the $s$-wave case, as originally proposed in [9]. Neglecting the real part of $f^{S}(p)$, the improved Jost function can be approximated as $\mathscr{J}_{\ell}^{\rm imp}(p)\simeq\alpha_{0}+i\gamma_{0}(p+p_{c})\simeq\mathscr{J}_{\ell}(p+p_{c})$, with $p_{c}$ being a regularising momentum given by $p_{c}={\rm Im}f_{\ell}^{S}(p)/\gamma_{0}^{2}=\mu(\sigma_{0}v)^{\rm ann,0}/4\pi\gamma_{0}^{2}$. Therefore, the improved Sommerfeld factor can also be approximated by the naive one evaluated at a momentum shifted by the critical value, $S^{\rm imp}_{0}(p)\simeq S_{0}(p+p_{c})$. In the Appendix C, we explicitly verify, that the improvement achieved by the RG method agrees well numerically with [9] and its extension [10].
@@
Fig. 2 shows the Sommerfeld factors as functions of momentum $p$ (normalized by the well width $R$) for the $s$-wave (left panel) and the $p$-wave (right panel). We set the mass parameter and the coupling constant to $\mu R=10$ and $\alpha_{D}=10^{-2}$, respectively. In both panels, the dashed and solid lines of the same color form a pair: the dashed line represents the conventional Sommerfeld factor $S_{\ell}(p)$ given in Eq. (2.25), while the corresponding solid line indicates the RG-improved Sommerfeld factor $S_{\ell}^{\rm imp}(p)$ in Eq. (3.24). The red solid line shows the ratio of the unitarity-bound annihilation cross section, which is given by the RHS of Eq. (3.3), to the tree-level annihilation cross section Eq. (3.36). For the $s$-wave case, we plot the Sommerfeld factors with the well depth chosen near the first zero-energy resonance, at $p_{V}=\pi/2-10^{-3}$ (purple), $p_{V}=\pi/2-10^{-5}$ (blue), and near the second zero-energy resonance, at $p_{V}=3\pi/2-10^{-5}$ (green). For the purple line, which is sufficiently far from the zero-energy resonance, the conventional Sommerfeld factor remains well below the unitarity bound and agrees well with the improved one. In contrast, for the blue and green lines, which are closer to the zero-energy resonance, both show unphysical behavior at small momenta, exceeding the unitarity bound. The RG-improved Sommerfeld factor resolves these issues and yields results consistent with unitarity. For the $p$-wave case, we plot the points $p_{V}=\pi-10^{-3}$ (purple) and $p_{V}=\pi-10^{-5}$ (blue) near the first zero-energy resonance, and $p_{V}=3\pi-10^{-5}$ (green) near the third zero-energy resonance. Unlike the $s$-wave case, the conventional Sommerfeld factors for all three parameter choices exhibit spike-like structures [8]. These arise from the fact that the $p$-wave can support resonant states, corresponding to zeros of the Jost function in the region of positive momentum. In particular, the blue and green lines exhibit resonant enhancements that exceed the unitarity bound. In contrast, the RG-improved Sommerfeld factor, shown as dashed lines, resolves such unphysical behavior: while it coincides with the conventional one for the purple line, which does not violate unitarity, the resonance peaks in the blue and green lines are properly regulated, yielding results consistent with unitarity.
@@ 4 Discussion
In this section, we provide further insights into the results obtained in the previous section by addressing two complementary aspects. First, we revisit the RG improvement procedure from the perspective of Wilsonian renormalization group [42], clarifying the validity and limitations of our method. This analysis helps identify conditions under which the RG improvement may fail or require modification. Second, as an advanced topic concerning the Sommerfeld effect, we briefly discuss the bound-state formation process [18, 43, 44], which can exhibit behaviors that violate the unitarity bound and therefore require careful treatment. These discussions aim to deepen the understanding of our formalism and its potential applicability, though they remain at a qualitative level and a more detailed treatment is left for future research.
@@ 4.1 Insights from Wilsonian renormalization group
where $G_{L}$ is a Green function associated with the long-range potential $V_{L}$. (For its explicit form, see [45]; however, the details are not essential for the following discussion.) The quantity $V_{S}(\Lambda)$ can be interpreted as the effective potential at the scale $p\sim\Lambda$, and the renormalization group equation describes its flow as a function of the momentum scale. The fixed points of the renormalization group equation (4.1) and their stability in the limit $p\to 0$ characterize the nature of the short-range contributions to the scattering process.
@@
where $\tilde{\delta}^{S}_{\ell}(p)$ is defined by the difference between the total phase shift $\delta_{\ell}(p)$ and that due to the long-range potential $\delta^{L}_{\ell}(p)$. $\mathcal{M}_{\ell}(p)$ is given by $\mathcal{M}_{\ell}(p)=\lim_{r\to 0}(d/dr)^{2\ell+1}((pr)^{\ell}\chi^{+}_{\ell,p}(r))/2^{\ell}\ell!\mathscr{J}_{\ell}(p)$, which is constructed using the Jost solution $\chi^{+}_{\ell,p}(r)$ (the solution of the Schrödinger equation (2.10) with boundary conditions imposed at infinity, $\chi^{+}_{\ell,p}(r)/h^{+}_{\ell}(pr)\to 1,r\to\infty$). The extra factors appearing in the modified effective range expansion Eq. (4.2) compared to the standard one Eq. (2.18) serve to remove the non-analytic momentum dependence arising from the long-range potential. This expansion has been used to subtract long-range effects such as the Coulomb force or one-pion exchange between nucleons [47, 29], thereby enabling the extraction of low-energy properties of the strong interaction. Therefore, if one allows the shape parameters $\tilde{a}_{\ell},\tilde{r}_{\ell}$ to take complex values to account for annihilation processes, Eq. (4.2) is expected to provide a useful parametrization of the self-scattering amplitude in situations where the short-range effect is nonperturbative, though it remains applicable even when the short-range effect is perturbative.
@@
Once the potential is regularized by some method and the Schrödinger equation is solved, the resulting $S$-matrix must reduce to the form of the modified effective range expansion, as it provides a general parametrization of the scattering amplitude for low-energy regime in the case where the potential comprises two separate contributions. In this case, the shape parameters become functions of the regularization parameters (and parameters contained in the long-range potential as well), but they are fixed to definite numerical values by matching to physical observables—namely, through renormalization. This approach is, in essence, equivalent to using the modified effective range expansion and fitting its parameters to scattering data at a suitable momentum scale. The key point, then, is whether the effect of the short-range potential is perturbative or nonperturbative. If the short-range interaction is perturbative, the corresponding contribution to the scattering amplitude can be computed directly from the UV theory using perturbation theory. In such cases, the RG–based treatment we have proposed is expected to provide a sufficiently accurate and much simpler approximation, particularly for computing unitarized Sommerfeld factors, compared to performing a matching based on the modified effective range expansion. Indeed, Ref. [9] discusses that, for the $s$-wave case, solving the Schrödinger equation with the short-range effect included yields results that agree well numerically with our expression given in Sec. 3.3 when the underlying short-range effect is perturbative, and we also confirm in Appendix C that the RG method and the direct solution of the Schrödinger equation in [9, 10] agree well numerically for the spherical well potential.
On the other hand, the analytical connection between the RG method and the formulations of Refs. [9] and [10] is conceptually subtle. Those works define short-range interactions—represented by delta functions and their derivatives—nonperturbatively and analyze the Sommerfeld effect, which leads them to encounter UV divergences immediately. By contrast, our framework defines the short-range interaction only at a perturbative level as in Eq. (2.23); as a result, the UV divergences appear sequentially with increasing perturbative order. Since in this work we restrict attention to the lowest-order approximation with UV-finite short-range input, we do not face those UV divergences. While extending our treatment to higher perturbative orders would indeed reintroduce UV issues, our view is that the essential unitarity question for the Sommerfeld effect is infrared (IR) rather than UV in nature; within that perspective, our method addresses unitarity correctly at the level considered. Establishing a more general analytical correspondence between the approaches based on exact solutions of the Schrödinger equation or on the modified effective range expansion and the RG-based framework—particularly in cases where the short-range interaction is perturbative for arbitrary partial waves—remains an important direction for future work.
@@ 4.2 Bound state formation
In the preceding discussion, we have considered only the simple process of direct annihilation. As a higher-order contribution to the annihilation process, it is known that slowly moving particles can form bound states through the emission of soft bosons that mediate their long-range interactions, which is referred to as bound state formation (BF) [18, 43, 44]. It has been pointed out that BF also exhibits resonance peaks at parameter values where the self-scattering cross section and the Sommerfeld factor resonate, in contradiction with the partial-wave unitarity bound [48, 49]. The unitarity violation problem in BF with unequal potentials in the initial and final states is discussed more generally in [50].
It would be interesting to apply the RG method to BF. Let us consider the simplest BF, where the mediator particle is a scalar boson $\varphi$, and that the transition from the scattering state $\mathcal{S}$ to a bound state $\mathcal{B}$ proceeds via the emission of a light boson with momentum $k$, mediated by a scalar transition operator $V_{k}$. At low energies, the initial state is dominated by the $s$-wave contribution and then, under the leading-order DWBA, the matrix element for this process is roughly given by
@@ 5 Conclusion
In the discussion, we examined our approach from the viewpoint of the Wilsonian renormalization group. From this perspective, our method under the assumption of a perturbative UV theory can be understood as a perturbative expansion around the Gaussian fixed point. The improvement of the amplitude through perturbative renormalization shares a parallel structure with the standard renormalized perturbation theory applied to weakly coupled field theories such as QED. Hence, the renormalization group approach is expected to offer a more tractable framework for unitarizing the Sommerfeld enhancement in systems with weak short-range interactions, compared to approaches based on exact solutions to the Schrödinger equation. Although the RG method shows good numerical agreement with that approach, establishing a precise analytical correspondence between the two frameworks for arbitrary partial waves remains an open problem.
@@ A.1 Riccati-Bessel functions
$j_{\ell}(z)=z^{\ell+1}\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!(2\ell+2n+1)!!}\left(\frac{z^{2}}{2}\right)^{n},\quad n_{\ell}(z)=\frac{1}{z^{\ell}}\sum^{\infty}_{n=0}\frac{(-1)^{n}(2\ell-2n-1)!!}{n!}\left(\frac{z^{2}}{2}\right)^{n}.$ (A.2)
@@
$h^{\pm}_{0}(z)=e^{\pm iz},$ $\quad h_{1}^{\pm}(z)=\left(1+\frac{i}{z}\right)e^{\pm i(z-\pi/2)}.$ (A.9)